Find the maximum value of z=6x−3x2 +2y





subject to the constraint





y − x2 = 2

1
Expert's answer
2022-05-17T11:13:17-0400

ANSWER The maximum value of "z=6x-3x^{2}+2y" subject to the constraint "y-x^{2}=2" is "z(3,11)=13" .

EXPLANATION

The set "\u0393" describing the constraint is "Γ=\left \{ \left ( x,2+x^{2} \right ): x\in \R \right \}" .On the set "\u0393" the function "z(x,y)=6x-3x^{2}+2y" has the form "z(x,2+x^{2})=6x-3x^{2}+2(2+x^{2})=-x^{2}+6x+4" for all "x\in\R" . Denote "h(x)=z(x,2+x^{2})." Therefore ,

"max_{Γ }z(x,y) =max_{\R}h(x)" .

Since "h'(x)=-2x+6" , then "h'(x)>0" if "x\in (-\infty,3)" and "h'(x)<0" if "x\in(3,+\infty)". "h'(3)=0" . Since "h" increases on the interval "(-\infty,3)" and decreases on the interval "(3,+\infty)" then "max_{(-\infty,3)}h(x)=h(3)=max_{(3,+\infty)}h(x)" . So "max_{\R}h(x)=h(3)=-9+18+4=13" "\Rightarrow max_{Γ}z(x,y)=13."